# TerraMind: Large-Scale Generative Multimodality for Earth Observation

### v3

@@ 4.3 Generation
After pretraining, TerraMind can generate tokens for any modality, conditioned on any subset of input modalities. The generative capabilities further unlock various zero-shot tasks, such as water body segmentation, land use segmentation, image captioning, and geolocation prediction. For the generation of image-like modalities, the decoder receives mask tokens for the modality to be generated and predicts the corresponding tokens based on the encoded input. For sequence-like modalities, the decoder generates the output autoregressively. After generating tokens from the target modality, the corresponding tokenizer decoder allows to map from token-space to image or text space.TerraMind further supports chained generation which ensures consistency across generated modalities. The chained generation represents a conditional probability distribution where the prior probability distribution $\mathbf{P(M_{1})}$ is given by the input modality $\mathbf{M_{1}}$, and all subsequent modalities $\mathbf{M_{2}}$ to $\mathbf{M_{K}}$ are generated conditioned on the input modality and potentially other generated modalities:
$P(\mathbf{M_{1}},\dots,\mathbf{M_{K}})=P(\mathbf{M_{1}})\prod_{i=2}^{K}P(\mathbf{M_{i}}\mid\mathbf{M_{1}},\dots,\mathbf{M_{i-1}}).$ (4)
@@ 4.4 Thinking-in-Modalities
Thinking in Modalities (TiM) is a recursive fine-tuning and inference technique designed to enhance multimodal learning by leveraging the generative capabilities of the model itself. Given an input $x\in\mathcal{X}$ (e.g., an optical satellite image), the model first generates additional synthetic modalities $\tilde{x}=f_{\text{gen}}(x)$ on a token-level using a learned generative function $f_{\text{gen}}$. These generated tokens are then concatenated with the original input and jointly processed by the downstream model $f$ (e.g., TerraMind encoder with a segmentation head), yielding the final output $y=f(x,f_{\text{gen}}(x))$. This formulation allows the model to reason over both observed and inferred modalities, effectively enriching the input space. TiM can leverage multiple generated modalities which are then generated in a chained approach (see Eq. 4). For example, for $k$ modalities, the input is augmented with newly generated modalities:
$\tilde{x}^{(k+1)}=\tilde{x}^{(k)}\cup f_{\text{gen}}(x^{(k)}),$ (5)
@@
This recursive augmentation mimics a chain-of-thought process, enabling the model to iteratively refine its internal representation and improve performance on downstream tasks, particularly in scenarios with missing modalities.
@@ 5 Experiments
In this section, we describe the performance gains resulting from TerraMind and experiment with the unlocked capabilities through the any-to-any generation of TerraMind. We especially investigate the quality of the artificially generated data as an input to subsequent finetuning tasks similar to chain–of-thought in LLMs.
@@ 5.1 Foundational experiments
S-1 61.01 63.94 (2.93$\uparrow$)
S-2 72.70 76.32 (3.62$\uparrow$)
LULC 71.77 70.96 (0.81$\downarrow$)
S-1 + S-2 73.83 76.74 (2.91$\uparrow$)
S-1 + LULC 73.86 73.76 (0.10$\downarrow$)
S-2 + LULC 75.65 77.04 (1.39$\uparrow$)
S-1 + S-2 + LULC 76.00 76.88 (0.88$\uparrow$)
Table 2: Task: Water body mapping on Sen1Floods11 [9], performance measured in IoU on water class. Model sizes and architectures are comparable. Left column: Late fusion of tokenizers. Effect of leveraging multimodal data in late-fusion fine-tuning of frozen modality-specific ViT-B models. The average improvement of full multimodality over the individual unimodal performance is 7.5pp IoU. Right column: Finetuning results of TerraMind.v1-B-single as a mid fusion approach based on masked correlation learning. Performance gains over late fusion are depicted as percentage points increases/decreases in parentheses.
Token-level fusion vs. late fusion. In Table 2 (right), we investigate the effects of fusing the inputs on a token level through masked token reconstruction compared to the late fusion from the previous paragraph. We observe that token-level fusion outperforms late fusion in water body mapping and the performance gains are particularly high when LULC data is not available. This suggests that early fusion captures an internal representation of the multimodal state—especially pronounced for LULC—that benefits fine-tuning. With those findings in mind, we will explore the effects of using additional multi-modal pixel-level input in a dual-scale pretraining in Section 5.5.
@@ 5.2 Generation experiments
(c) Input: S-2 L2A data capturing Nothern Spain in January 2025.
@@
Figure 6: Generated Sentinel-1 data using TerraMind over Singapore and Northern Spain.
@@
TerraMind supports the generation of any modality from any other modality. In the following, we provide examples of the generation performance starting from: (i) an information-rich modality, like optical S-2 L2A data, and (ii) minimal information based on the geolocation. In Figure 3, we observe that TerraMind performs very well in generating image-like modalities like radar data, land use maps, and digital elevation maps from optical S-2 L2A data. We provide a quantitative overview on the quality of the generations on unseen validation data in Table 3. Overall, we observe an interesting asymmetry in the generative performance of TerraMind where (a) radar-to-optical generation achieves reasonable generation quality in terms of SSIM and PSNR – indicating structural and visual fidelity with some perceptual degradation – and (b) optical-to-radar generation yields higher PSNR values but lower SSIM, suggesting visually plausible outputs that lack strong structural alignment. The quality of generated DEM suggest to be structurally very strong, but noisy. The errors for DEM generations suggest that the generations structurally make sense but the level of altitude is difficult to infer for the model. We compare these scores with the reconstruction quality of the auto-encoding tokenizers in the supplementary material that can serve as upper bounds. Additionally, we provide experiments on the generation quality using token-level instead of pixel-level inputs in the supplementary material. Finally, we demonstrate the quality of large-scale generations across Singapore and Northern Spain in Figures 19 and 21.
@@ 5.3 Zero-shot experiments
Water body mapping. In Table 4, we compare the zero-shot performance of TerraMind with its fine-tuned performance and other finetuned benchmarks for water body mapping. Overall, TerraMindv1-B-single achieves a zero-shot IoU of 45.4% compared to SOTA-level fine-tuning performance of 84.4% of DeCUR. In ablations with TerraMindv1-single trained on DynamicWorld LULC data, we boost this to up to 69.8% suggesting that TerraMind harnesses up to over 80% of the SOTA performance in zero-shot setting. Additionally, it’s notable that none of the benchmarking model can be applied in a zero-shot context, highlighting the relevance of TerraMind’s capabilities.
@@
Figure 7: Prediction distribution of the land use class “bare land” with a sampling temperature of $T=1.0$. TerraMind has an accurate internal representation of the geolocation of specific contexts, like land use classes.
@@ 5.4 Few-shot experiments
TerraMind is pretrained on a cross-modal patch classification objective. Therefore, we expect a well-structured latent space that clusters different concepts accurately. To investigate our hypothesis, we apply 1-Nearest-Neighbor (1-NN) classification experiments in the community-standard setting of 1-shot 5-way on two datasets: EuroSAT and METER-ML. In those experiments, there are no weight updates of any kind, so that we can assess the quality of the embedding space structure.
@@
In Table 5, we observe that TerraMind outperforms several other benchmarks from both the CV and EO domain on the EuroSAT dataset by at least 10pp in accuracy. Our results further show that for methane source classification on METER-ML, TerraMind outperforms benchmark models and generalizes to high-resolution NAIP data with one order of magnitude higher resolution than the pre-training data. We present additional experiments with other few-shot settings in the supplementary material.
@@ 5.5 Fine-tuning experiments
Model BurnSr MADOS PASTIS Sen1Fl11 FBP DEN CTM-SS SN7 AI4Farms Avg. mIoU Avg. Rank CROMA 82.42 67.55 32.32 90.89 51.83 38.29 49.38 59.28 25.65 55.29 6.61 DOFA 80.63 59.58 30.02 89.37 43.18 39.29 51.33 61.84 27.07 53.59 8.22 GFM-Swin 76.90 64.71 21.24 72.60 67.18 34.09 46.98 60.89 27.19 52.42 10.00 Prithvi 1.0 100M 83.62 49.98 33.93 90.37 46.81 27.86 43.07 56.54 26.86 51.00 11.00 RemoteCLIP 76.59 60.00 18.23 74.26 69.19 31.78 52.05 57.76 25.12 51.66 11.22 SatlasNet 79.96 55.86 17.51 90.30 50.97 36.31 46.97 61.88 25.13 51.65 10.67 Scale-MAE 76.68 57.32 24.55 74.13 67.19 35.11 25.42 62.96 21.47 49.43 11.44 SpectralGPT 80.47 57.99 35.44 89.07 33.42 37.85 46.95 58.86 26.75 51.87 10.11 S.-S12-MoCo 81.58 51.76 34.49 89.26 53.02 35.44 48.58 57.64 25.38 53.02 10.06 S.-S12-DINO 81.72 49.37 36.18 88.61 51.15 34.81 48.66 56.47 25.62 52.51 10.89 S.-S12-MAE 81.91 49.90 32.03 87.79 51.92 34.08 45.80 57.13 24.69 51.69 12.39 S.-S12-Data2Vec 81.91 44.36 34.32 88.15 48.82 35.90 54.03 58.23 24.23 52.22 10.72 UNet Baseline 84.51 54.79 31.60 91.42 60.47 39.46 47.57 62.09 46.34 57.58 4.89 ViT Baseline 81.58 48.19 38.53 87.66 59.32 36.83 44.08 52.57 38.37 54.13 10.28 TerraMindv1-B-single 84.00 65.01 40.80 90.32 52.66 59.71 27.71 TerraMindv1-B 82.42 69.52 40.51 90.62 59.72 37.87 55.80 60.61 28.12 58.35 3.94 TerraMindv1-L 82.93 75.57 43.13 90.78 63.38 37.89 55.04 59.98 27.47 59.57 3.44
@@
Table 6: Performance evaluation of TerraMind across nine benchmark datasets using the PANGAEA evaluation protocol. Higher mIoU values ($\uparrow$) and lower rank values ($\downarrow$) are reported. The best model per column is highlighted in bold and the second best is underscored. We indicate unimodal datasets with , demonstrating that TerraMind outperforms other benchmarks in both unimodal and multimodal settings. All foundation model encoders are frozen, while U-Net and ViT baselines are trained from scratch for each specific task.
@@
Besides the novel capabilities that TerraMind introduces, we benchmark the fine-tuning performance of TerraMind in both unimodal and multimodal settings following the community-standard PANGAEA benchmark [49]. We summarize the results in Table 6. Overall, TerraMindv1-B outperforms all other GeoFMs by at least 3pp avg. mIoU. Importantly, we observe that TerraMind is the only foundation model approach in EO that across the PANGAEA benchmark outperforms task-specific U-Net models. Performance is approximately 2pp avg. mIoU higher for TerraMindv1-L, with a peak of 5pp in multimodal datasets. Furthermore, TerraMindv1-L outperforms also specialised ViT baselines by 5pp avg. mIoU. Finally, TerraMindv1-B outperforms TerraMindv1-B-single by 1pp. avg. mIoU. The latter result demonstrates the benefit of using dual-scale features when comparing v1-B-single with v1-B. Note that per suggestion of the PANGAEA authors, we exclude the xView2 and BioMassters task as we could not reproduce the reported performances. Finally, we assess the impact of leveraging multimodal data as input to TerraMindv1-B compared to utilizing either optical or radar data as unimodal input to better understand the effect of leveraging multimodal data in finetuning. We observe that across all three multimodal tasks, TerraMindv1-B performs best with access to both optical and radar data with a significant improvement for CTM-SS.
PASTIS Sen1Fl11 CTM-SS
@@
Multimodal 40.51 90.62 55.80
Table 7: Benefit of using multimodal input in TerraMind.v1-B on multimodal datasets of PANGAEA. We report the mIoU (%) $\uparrow$.
@@ 5.6 Thinking in modalities
Based on the generative capabilities of TerraMind and the fact that multi-modal data improved downstream task performances in many scenarios, we additionally investigate the value of artificially generated data for finetuning tasks. Borrowing on the idea of chain-of-thought finetuning of large language models, we use TerraMind to generate artificial data as an intermediate step to improve finetuning. We provide an overview on this new method in the beginning of this paper in Figure 1.
@@
We employ the TiM tuning approach on water body mapping and during finetuning use radar or optical data to generate additional artificial LULC data which we leverage along the input. Our results in Table 8 indicate a superior performance of the TiM finetuning compared to leveraging uni-modal data by up to 2pp mIoU. We also note that TiM tuning strongly benefits from a trainable encoder. This finding points us in the direction of TerraMind being able to generate data that improve finetuning to downstream applications.
@@ 6 Conclusion
TerraMind’s pre-training approach of combining token-level and pixel-level information has unlocked a range of new model capabilities in Earth Observation. TerraMind demonstrates not only beyond state-of-the-art performance when tested against 12 other geospatial models on the PANGEA benchmark, it also represents the first truly generative multimodal model in the domain. Because of the ability of integrating heterogeneous and diverse data sources demonstrated by the Terramind’s approach, we expect that TerraMind-like models will expand soon to multi-temporal, multi-resolution, and hyperspectral data. This will allow to maximise impact in real world applications and fully leverage the data rich ecosystem available in the Earth Observation domain.
@@ 7 TerraMesh Dataset
Figure 8: Visualization of the spatial-temporal alignment across modalities in TerraMesh. S-2 L2A uses IRRG pseudo-coloring and S-1 RTC is visualized in db scale as VH-VV-VV/VH. Copernicus DEM is scaled based on the image value range.
TerraMind.v1-B-single was pre-trained on a subset of TerraMesh with one million samples, specifically the SSL4EOS12 v1.1 locations, using only four image modalities: S-2 L2A, S-1 GRD, DEM, and LULC. Additionally, we performed continuous pre-training with image captions. These captions were created using LLaVA-Next [37] and Overture Maps data. The automated captioning pipeline includes a prompt with a chain-of-thought process to generate diverse captions. The captioning model is asked to generate three question-answer pairs and describe the full image later. We use the S-2 RGB bands and Overture base layer tags as inputs. Domain experts evaluated a subset of 1.3k captions, resulting in 69% of the captions without any hallucinations while the average completeness scores were 3.87 on a scale from 0 to 5. The captions will be released as an additional preprint during the review process of this paper.

### v4

@@ 4.3 Generation
Once pretrained, TerraMind can generate tokens for any modality, conditioned on any subset of input modalities. The generative capabilities unlock various zero-shot tasks, such as water body segmentation. For the generation of image-like modalities, the decoder receives mask tokens for the modality to be generated and predicts the corresponding tokens based on the encoded input. For sequence-like modalities, the decoder generates the output autoregressively. After generating tokens from the target modality, the corresponding tokenizer decoder allows to map from token-space to image or text space. TerraMind further supports chained generation which ensures consistency across generated modalities. The chained generation represents a conditional probability distribution where the prior probability distribution is determined by the input modality, and all subsequent modalities are generated conditioned on the input modality and potentially other generated modalities.
@@ 4.4 Thinking-in-Modalities
Thinking in Modalities (TiM) is a recursive fine-tuning and inference technique designed to enhance multimodal learning by leveraging the generative capabilities of the model itself. Given an input $x\in\mathcal{X}$ (e.g., an optical satellite image), the model first generates additional synthetic modalities $\tilde{x}=f_{\text{gen}}(x)$ on a token-level using a learned generative function $f_{\text{gen}}$. These generated tokens are then concatenated with the original input and jointly processed by the downstream model $f$ (e.g., TerraMind encoder with a segmentation head), yielding the final output $y=f(x,f_{\text{gen}}(x))$. This formulation allows the model to reason over both observed and inferred modalities, effectively enriching the input space. TiM can leverage multiple generated modalities which are then generated in a chained approach. For example, for $k$ modalities, the input is augmented with newly generated modalities:
$\tilde{x}^{(k+1)}=\tilde{x}^{(k)}\cup f_{\text{gen}}(\tilde{x}^{(k)}),$ (2)
@@
This recursive augmentation mimics a chain-of-thought process, enabling the model to iteratively refine its internal representation, particularly in scenarios with missing modalities.
@@ 5 Experiments
In this section, we describe the performance gains resulting from TerraMind and experiment with the unlocked capabilities of any-to-any generation and Thinking-in-Modalities.
@@ 5.1 Foundational experiments
[…]
Table 2: Water body mapping on Sen1Floods11 [9] measured in IoU on water class. Model sizes and architectures are comparable. Left column: Late fusion of tokenizers. The average improvement of full multimodality over the individual unimodal performance is 7.5pp IoU. Right column: Finetuning results of TerraMindv1-B-single as a mid fusion approach based on masked correlation learning. Gains over late fusion in percentage points in parentheses.
Token-level fusion vs. late fusion. In Table 2 (right), we investigate the effects of fusing the inputs on a token level through masked token reconstruction. We observe that token-level fusion outperforms late fusion. The performance gains are particularly high when LULC data is not available. This suggests that early fusion captures an internal representation of the multimodal state—especially pronounced for LULC—that benefits fine-tuning. With those findings in mind, we will explore the effects of using additional multi-modal pixel-level input in a dual-scale pretraining in Section 5.5.
@@ 5.2 Generation experiments
TerraMind supports any-to-any generation. In the following, we provide examples of the generation performance starting from: (i) an information-rich modality, like optical S-2 L2A data, and (ii) minimal information based on the geolocation. In Figure 3, we observe that TerraMind performs strongly in generating image-like modalities like S-1, LULC, and DEM from optical S-2 L2A data. We provide a quantitative overview on the quality of the generations on unseen validation data in Table 3. Overall, we observe an interesting asymmetry in the generative performance of TerraMind where (a) radar-to-optical generation achieves reasonable quality in terms of SSIM and PSNR – indicating structural and visual fidelity with some perceptual degradation – and (b) optical-to-radar generation yields higher PSNR values but lower SSIM, suggesting visually plausible outputs that lack strong structural alignment. The quality of generated DEM suggests to be structurally very strong, but noisy. The errors for DEM generations suggest that the level of altitude is difficult to infer for the model. We compare these scores with the reconstruction quality of the auto-encoding tokenizers in the supplementary material that can serve as upper bounds. Additionally, we provide experiments on the generation quality using token-level instead of pixel-level inputs. Finally, we demonstrate the quality of generations at kilometer scale in Figures 19 and 21.
@@
(c) Input: S-2 L2A data capturing Northern Spain in January 2025.
@@
Figure 6: Generated S-1 imagery using TerraMind. We provide large-scale visualizations in the supplementary material.
@@ 5.3 Zero-shot experiments
Water body mapping. In Table 4, we compare the zero-shot performance of TerraMind with its fine-tuned performance and other finetuned benchmarks for water body mapping. Overall, TerraMindv1-B achieves a zero-shot IoU of 45.4% compared to SOTA-level fine-tuning performance of 82.2% of DeCUR. In ablations with TerraMindv1-B-single trained on DynamicWorld LULC data, we boost this to up to 69.8% suggesting that TerraMind harnesses up to over 80% of the SOTA performance in zero-shot setting. Additionally, it’s notable that none of the benchmarking model can be applied in a zero-shot context, highlighting the relevance of TerraMind’s capabilities.
@@
Figure 7: Prediction distribution of the land use class “bare land” with a sampling temperature of $T=1.0$ using TerraMindv1-B-single. TerraMind has an accurate internal representation of the geolocation of specific contexts, like land use classes.
@@ 5.4 Few-shot experiments
TerraMind is trained via a cross-modal patch classification objective. Thus, we expect a well-structured latent space that clusters different concepts accurately. To investigate our hypothesis, we apply 1-Nearest-Neighbor (1-NN) classification experiments in the community-standard setting of 1-shot 5-way on two datasets: EuroSAT and METER-ML. In those experiments, there are no weight updates of any kind, so that we can assess the quality of the embedding space structure. In Table 5, we observe that TerraMind outperforms several other benchmarks from both the CV and EO domain on the EuroSAT dataset by at least 10pp in accuracy. Our results further show that for methane source classification on METER-ML, TerraMind outperforms benchmark models and generalizes to high-resolution NAIP data with one order of magnitude higher resolution than the pre-training data. We present additional experiments with other few-shot settings in the supplementary material.
@@ 5.5 Fine-tuning experiments
Model BurnSr MADOS PASTIS Sen1Fl11 FBP DEN CTM-$S$ SN7 AI4Farms Avg. mIoU Avg. Rank CROMA 82.42 67.55 32.32 90.89 51.83 38.29 49.38 59.28 25.65 55.29 6.61 DOFA 80.63 59.58 30.02 89.37 43.18 39.29 51.33 61.84 27.07 53.59 8.22 GFM-Swin 76.90 64.71 21.24 72.60 67.18 34.09 46.98 60.89 27.19 52.42 10.00 Prithvi 1.0 100M 83.62 49.98 33.93 90.37 46.81 27.86 43.07 56.54 26.86 51.00 11.00 RemoteCLIP 76.59 60.00 18.23 74.26 69.19 31.78 52.05 57.76 25.12 51.66 11.22 SatlasNet 79.96 55.86 17.51 90.30 50.97 36.31 46.97 61.88 25.13 51.65 10.67 Scale-MAE 76.68 57.32 24.55 74.13 67.19 35.11 25.42 62.96 21.47 49.43 11.44 SpectralGPT 80.47 57.99 35.44 89.07 33.42 37.85 46.95 58.86 26.75 51.87 10.11 S.-S12-MoCo 81.58 51.76 34.49 89.26 53.02 35.44 48.58 57.64 25.38 53.02 10.06 S.-S12-DINO 81.72 49.37 36.18 88.61 51.15 34.81 48.66 56.47 25.62 52.51 10.89 S.-S12-MAE 81.91 49.90 32.03 87.79 51.92 34.08 45.80 57.13 24.69 51.69 12.39 S.-S12-Data2Vec 81.91 44.36 34.32 88.15 48.82 35.90 54.03 58.23 24.23 52.22 10.72 UNet Baseline 84.51 54.79 31.60 91.42 60.47 39.46 47.57 62.09 46.34 57.58 4.89 ViT Baseline 81.58 48.19 38.53 87.66 59.32 36.83 44.08 52.57 38.37 54.13 10.28 TerraMindv1-B 82.42 69.52 40.51 90.62 59.72 37.87 55.80 60.61 28.12 58.35 3.94 TerraMindv1-L 82.93 75.57 43.13 90.78 63.38 37.89 55.04 59.98 27.47 59.57 3.44
@@
Table 6: Performance evaluation of TerraMind using the PANGAEA evaluation protocol indicates higher mIoU values ($\uparrow$) and lower rank values ($\downarrow$). The best model per column is highlighted in bold, the second best is underscored. We indicate unimodal datasets with . Encoders are frozen for pretrained models, while U-Net and ViT baselines are trained from scratch for each specific task.
@@
Besides the novel capabilities that TerraMind introduces, we benchmark the fine-tuning performance of TerraMind in both unimodal and multimodal settings following the community-standard PANGAEA benchmark [49]. We summarize the results in Table 6. Overall, TerraMindv1-B outperforms all other GeoFMs by at least 3pp avg. mIoU. Importantly, we observe that TerraMind is the only foundation model approach in EO that across the PANGAEA benchmark outperforms task-specific U-Net models. Performance increases by approximately 2pp avg. mIoU for TerraMindv1-L, with a peak of 5pp in multimodal datasets. Furthermore, TerraMindv1-L outperforms also specialised ViT baselines by 5pp avg. mIoU. Note that per suggestion of the PANGAEA authors, we exclude the xView2 and BioMassters task as we could not reproduce the reported performances. Finally, we assess the impact of leveraging multimodal data as input to TerraMindv1-B compared to utilizing either optical or radar data as unimodal input to better understand the effect of leveraging multimodal data in finetuning. We observe that across all three multimodal tasks, TerraMindv1-B performs best with access to both optical and radar data.
PASTIS Sen1Fl11 CTM-$S$
@@
S-1 + S-2 40.51 90.62 55.80
Table 7: Benefit of using multimodal input in the PANGAEA benchmark reported in mIoU (%) $\uparrow$.
@@ 5.6 Thinking in modalities
We additionally evaluate the value of TiM tuning on water body mapping. We use S-1 or S-2 to generate artificial LULC data as additional input. Our results in Table 8 indicate a superior performance of TiM tuning compared to leveraging uni-modal data by up to 2pp mIoU. This finding points us in the direction of TerraMind being able to generate data that improve downstream task performance. We provide additional results in the appendix.
@@ 6 Conclusion
TerraMind’s approach of combining token-level and pixel-level data has unlocked a range of new model capabilities in EO. TerraMind demonstrates not only beyond state-of-the-art performance in community-standard benchmarks, it also represents the first fully generative multimodal model in the domain. Because of the ability of integrating heterogeneous data sources, we expect that TerraMind-like models will expand to multi-temporal, multi-resolution, and hyperspectral data to fully leverage the data rich ecosystem available in the Earth Observation domain.
@@ 7 TerraMesh Dataset
Figure 8: Visualization of the spatial-temporal alignment across modalities in TerraMesh. S-2 L2A uses IRRG pseudo-coloring and S-1 RTC is visualized in db scale as VH-$V$-VV/VH. Copernicus DEM is scaled based on the image value range.
TerraMind.v1-B-single was pre-trained on a subset of TerraMesh with one million samples, specifically the SSL4EOS12 v1.1 locations, using only four image modalities: S-2 L2A, S-1 GRD, DEM, and LULC. Additionally, we performed continuous pre-training with image captions. These captions were created using LLaVA-Next [37] and Overture Maps data [47]. The automated captioning pipeline includes a prompt with a chain-of-thought process to generate diverse captions. The captioning model is asked to generate three question-answer pairs and describe the full image later. We use the S-2 RGB bands and Overture base layer tags as inputs. Domain experts evaluated a subset of 1.3k captions, resulting in 69% of the captions without any hallucinations while the average completeness scores were 3.87 on a scale from 0 to 5.